The total cost of 8 gallons of regular unleaded gasoline and 12 gallons of premium unleaded is $27.84. premium unleaded gasoline costs $0.17 more per gallon than regular unleaded. Find the price per gallon for each grade of gasoline. Please show how to set this equation up and answer it.
Answer by LilSkittleMd(119)

Let r=regular, p=premium
First set up the two equations
8r+12p=27.84
p=r+0.17
Substitute for p
8r+12(r+0.17)=27.84
Distribute
8r+12r+2.04=27.84
Combine like terms
20r+2.04=27.84
Subtract 2.04 from both sides of the equation
20r=25.80
Divide both sides of the equation by 20
r=1.29
Now plug in the value for r into one of the original equations
p=1.29+0.17
p=1.46
One gallon of regular unleaded gasoline is $1.29
One gallon of premium unleaded gasoline is $1.46